OpenStudy (anonymous):

What is the slope of the line below? If necessary, enter your answer as a fraction in lowest terms, using the slash ( / ) as the fraction bar. Do not enter your answer as a decimal number or an equation. (2,2) (-1,-4)

myininaya (myininaya):

subtract y in the first ordered pair from the y in the second ordered pair this difference will be your rise subtract x in the first ordered pair from the x in the second ordered pair this difference will be your run

myininaya (myininaya):

If you are more likely to understand formulas the formula is: \[\text{ slope } =\frac{ y_2-y_1}{x_2-x_1} \text{ where the points the line go through are } \\ (x_1,y_1) \text{ and } (x_2,y_2)\]

OpenStudy (anonymous):

2-(-1)= 3 2-(-4)=6 Answer (3,6)?